OpenStudy (anonymous):

How do you simplify 4(Q/2)^1/2

OpenStudy (anonymous):

The answer is supposed to be √8Q but I don't know how they got that...

OpenStudy (anonymous):

\[4\sqrt{Q/2}\]you can't include 4 as part of the radicand since there is a sign for grouping. therefor, 4 should be multiplied to the sqrt of Q/2

jhonyy9 (jhonyy9):

this is 4sqrt(Q/2)=sqrt(16Q/2)=sqrt(8Q)=sqrt(4*2Q)=2sqrt(2Q)